# Almost Complete Intersections: Regularity of powers and Rees algebra

### v1

@@ 2. Regularity of powers of Almost Complete Intersections
    For an odd integer $t\geq 5$, $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ is a $d$-sequence generating set of an almost complete intersection ideal of grade $3$ and type $t-3$.
@@
    For an even integer $t\geq 6$, $\{\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X),\operatorname{Pf}(X)\}$ is a $d$-sequence generating set of an almost complete intersection ideal of grade $3$ and type $t-3$.
@@
We have $\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X)\}$ forms a regular sequence from [7, C.2 Lemma]. Further, since $J$ is unmixed, from [13, Proposition 1.3], $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ forms a $d$-sequence.
Similarly $(2)$ can be seen as a consequence of [7, C.6 Lemma] and [13, Proposition 1.3]. ∎
@@
Assume the notation and hypotheses as in Setup 3. Then for $m\geq 2$,
@@
    If $t\geq 5$ is an odd integer such that , then for $i=1,2,3$,
    $\operatorname{reg}\left(\dfrac{B}{\langle f_{1},\ldots,f_{i}\rangle+I^{m}}\right)=\left(\frac{t-1}{2}\right)(m+2)-5.$
@@
    $\operatorname{reg}\left(\dfrac{B}{\langle f_{1},\ldots,f_{i}\rangle+I^{m}}\right)=\left(\frac{t}{2}\right)(m+2)-7.$
@@
The result can be seen as a consequence of Theorem 2.3 and Lemma 2.5
@@ 3. Rees algebra of Almost Complete Intersections
This section discusses the resolutions of the Rees algebra of some classes of ideals $I$ of linear type satisfying a relation between their minimal number of generators and their respective heights. The resolution is obtained as the mapping cone between the resolution of $I$ and the Buchsbaum-Rim complex, coming corresponding to the generators of $I$. The idea has been discussed in [5] for constructing resolutions of equigenerated ideals. We try to extend the result to include the non-equigenerated case as well. As a consequence of this, minimal bigarded free resolutions of almost complete intersections are obtained.
The following are the building blocks for constructing the minimal bigraded free resolutions given in Proposition 3.1.
@@
    $\epsilon(e_{i})=\sum_{m\subset k,\,|m|=2}\text{sgn}(m\subset k)(\text{det}(\psi_{m}))e_{k\backslash m}$
@@
    $\sigma_{i}(\beta\otimes\gamma)=\sum_{s,t}[\psi^{*}(\beta_{s}^{\prime})](\gamma_{t}^{\prime})\cdot\gamma_{t}^{\prime\prime}\otimes\beta_{s}^{\prime\prime}.$
@@
Next we mention a map of complexes from $\mathbb{C.}$ to $\mathbb{F.}$ mentioned in [5].
Let $\nu$ be a map of complexes from the Koszul complex corresponding to $\phi_{1}$, denoted by (.$(\wedge^{.}A^{l},c.)$, to the Buchsbaum-Rim complex (.$(\mathbb{C.},\sigma.)$, obtained by extending the identity map at degree 00 and degree $1$ such that the diagram commutes. This map $\nu$ helps in defining the required map of complexes $\alpha:\mathbb{C.}\rightarrow\mathbb{F.}$ which has the following form.
For $21\leq i\leq l-$ 2, $\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\rightarrow F_{i+1}\otimes S$ is given as the composition of the following two maps.
$\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\stackrel{{\scriptstyle p\otimes\tau}}{{\longrightarrow}}\wedge^{1}F\otimes\wedge^{i+1}F\stackrel{{\scriptstyle\tiny{\begin{bmatrix}y_{1}&\cdots&y_{l}\end{bmatrix}}\otimes\nu_{i+1}}}{{\longrightarrow}}S\otimes F_{i+1}$ (5)
@@
$\cdots\stackrel{{\scriptstyle\delta_{4}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 3}S(-j,-1)^{\beta_{4j}}\\ \oplus\\ C_{2}\end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 2}S(-j,-1)^{\beta_{3j}}\\ \oplus\\ C_{1}\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 1}S(-j,-1)^{\beta_{2j}}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0.$ (6)
@@
$C_{i}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$
are the modules which occur in the Buchsbaum-Rim complex $\mathbb{C}.$ corresponding to the matrix $\begin{bmatrix}f_{1}&\cdots&f_{n+1}\\ y_{1}&\cdots&y_{n+1}\\ \end{bmatrix}$ with differentials denoted by $\sigma_{i}$. Let $\alpha:\mathbb{C}.\rightarrow\mathbb{F}.$ be the map of complexes given in Equation (5). Then for all $i=1,\ldots,\operatorname{proj\,dim}_{S}\mathcal{R}(I)$, the differentials $\delta_{i}$ are of the following form.
$\delta_{1}=\begin{bmatrix}y_{1}&\ldots&y_{n+1}\end{bmatrix}\circ\phi_{2},\;\;\delta_{2}=\begin{pmatrix}\alpha_{1}&\phi_{3}\end{pmatrix},\;\;\delta_{i}=\begin{pmatrix}\phi_{i+1}&\alpha_{i-1}\\ O&\sigma_{i-1}\end{pmatrix}.$
@@
Here for $i=1,\ldots,n+1$, $\psi^{*}(g_{1}^{*})(e_{i})=f_{i}$ and $\psi^{*}(g_{2}^{*})(e_{i})=y_{i}$ which determines the entries in the matrices corresponding to the differentials $\sigma_{i}^{\prime}s$. Then the way the differentials are defined in Complex (8), one obtains
$C_{i}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$
for $i=1,\ldots,n-1$.
Consider (.$(\wedge^{.}B^{n+1},k.)$ to be the Koszul complex corresponding to $\phi_{1}$. Now since for $\{j_{1},j_{2}\}\subseteq[n+1]$ such that $j_{1}\neq j_{2}$, $k_{2}(e_{j_{1}}\wedge e_{j_{2}})=f_{j_{1}}e_{j_{2}}-f_{j_{2}}e_{j_{1}}\in$ Ker ϕ1=subscriptitalic-$\phi_{1}=$ Im $\phi_{2}$, there exists $\beta_{j_{1}j_{2}}\in F_{2}$ such that $\phi_{2}(\beta_{j_{1}j_{2}})=i(k_{2}(e_{j_{1}}\wedge e_{j_{2}}))$ where $i$ is the natural map from $\wedge^{0}B^{n+1}$ to $F_{1}$. Let $\alpha:\mathbb{C.}\rightarrow\mathbb{F.}$ be the map of complexes. Then $\alpha_{1}:\wedge^{3}{S^{n+1}}\otimes(\operatorname{Sym}_{0}S^{2})^{*}\rightarrow S\otimes F_{2}$ will be defined as, for $\{j_{1},j_{2},j_{3}\}\subseteq[n+1]$, $\alpha_{1}((e_{j_{1}}\wedge e_{j_{2}}\wedge e_{j_{3}})\otimes 1)=y_{j_{1}}\beta_{j_{2}j_{3}}-y_{j_{2}}\beta_{j_{1}j_{3}}+y_{j_{3}}\beta_{j_{1}j_{2}}.$
With the details above, one obtains the required resolution (6) by considering the mapping cone of $\alpha$ (arguments similar to the ones in [5]). Further, the resolution is minimal since the entries in the matrices given by the differentials $\delta_{i}$ are from $\mathfrak{m}S$, where $\mathfrak{m}$ is the bigraded maximal ideal of $S$. Note that the linear shifts in the second variable is contributed by the differentials in the minimal free resolution of $B/I$ and the higher degree shifts for the same are contributed by $\alpha_{i}^{\prime}s$ and $\sigma_{i}^{\prime}s$. ∎
@@
Note that in Proposition 3.1 if $I$ is considered to be a graded ideal of $B$ with $\mu(I)=\text{ht}(I)+1$ and not necessarily of linear type, then the Complex (6) gives a minimal bigraded free resolution of the $\operatorname{Sym}(I)$.
@@
Following observations can be made on the nature of the generators of $I$ as a consequence of the explicit minimal bigraded free resolutions of the symmetric algebra of $I$ obtained from Proposition 3.1.
Note that for an ideal $I$ in $B$, it’s generating set $\{a_{1},\ldots,a_{n}\}$ forms a strong $s$-sequence (w.r.t an admissible term order for the monomials in $y_{i}$ with $y_{1}<y_{2}<\ldots<y_{n}$) if for $\operatorname{Sym}(I)\cong S/J$ where $S$ is a bigraded polynomial ring $S=B[y_{1},\ldots,y_{n}]$ and $J$ the defining ideal of the symmetric algebra, $in(J)=\langle L_{1}y_{1},\ldots,L_{n}y_{n}\rangle$ where $L_{i}=(a_{1},\ldots,a_{i-1}:a_{i})$ and $L_{1}\subseteq L_{2}\subseteq\ldots\subseteq L_{n}$ (cf. [16]).
@@
From the minimal bigraded free resolution of the $\operatorname{Sym}(I)$ in Complex (6), it is clear that the $y$-regularity of the $\operatorname{Sym}(I)$ is 00. The result then follows from [30, Corollary 3.2].
Further when the ideal is of linear type, $I$ is generated by a $d$-sequence can be seen as a consequence of [25, Theorem 2.9]. ∎
@@
    $d$-sequence ideals satisfying the condition that the grade of the ideal is one more than the minimal number of generators of the ideal.
  3. (3)
    Almost complete intersections.
In the remaining part of this section, we give the explicit resolutions of the Rees algebra of some classes of ideals satisfying the hypothesis of Proposition 3.1, in particular almost complete intersections of grade $2$, grade $3$ etc.

### v2

@@ 2. Regularity of powers of Almost Complete Intersections
    For an odd integer $t\geq 5$, $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ is a $d$-sequence such that $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X)\}$ forms a regular sequence in $K[X]$.
@@
    For an even integer $t\geq 6$, $\{\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X),\operatorname{Pf}(X)\}$ is a $d$-sequence such that $\{\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X)\}$ forms a regular sequence in $K[X]$.
@@
We have $\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X)\}$ forms a regular sequence from [8, C.2 Lemma]. Further, since $J$ is unmixed, from [14, Proposition 1.3], $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ forms a $d$-sequence.
Similarly $(2)$ can be seen as a consequence of [8, C.6 Lemma] and [14, Proposition 1.3]. ∎
@@
Assume the notation and hypotheses as in Setup 2. Then for $m\geq 2$,
@@
    If $t\geq 5$ is an odd integer, then for $i=1,2,3$,
    $\operatorname{reg}\left(\dfrac{B}{I^{m}}\right)=\left(\frac{t-1}{2}\right)(m+2)-5.$
@@
    $\operatorname{reg}\left(\dfrac{B}{I^{m}}\right)=\left(\frac{t}{2}\right)(m+2)-7.$
@@
We have $I$ to be an almost complete intersection ideal of $B=K[X]$, a standard graded polynomial ring with $X$ being a skew-symmetric matrix of the form (5), of grade $3$ and type $t-3$ for $t\in\mathbb{N}$ and $t\geq 5$.
  1. (1)
    If $t$ is an odd integer, then from Equation (6), $I$ is generated by the Pfaffians $\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)$ which are of degrees $\frac{t-3}{2},\frac{t-1}{2},\frac{t-1}{2},\frac{t-1}{2}$, respectively. From Remark 2.6(2), we have $\text{reg}(B/I)=t-4<\left(\frac{t-3}{2}\right)+\left(\frac{t-1}{2}\right)+\left(\frac{t-1}{2}\right)-4+1$. Thus, the result can be seen as a consequence of Lemma 2.7, Note 2.5 and the simplification of the formula derived from Theorem 2.3.
  2. (2)
    If $t$ is an even integer, then from Equation (7), $I$ is generated by the Pfaffians $\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X),\operatorname{Pf}(X)$ which are of degrees $\frac{t-2}{2},\frac{t-2}{2},\frac{t-2}{2},\frac{t}{2}$, respectively. From Remark 2.6(2), since $\text{reg}(B/I)=t-4<\left(\frac{t-2}{2}\right)+\left(\frac{t-2}{2}\right)+\left(\frac{t-2}{2}\right)-4+1$, similar to the previous case, the result follows from Lemma 2.7, Note 2.5 and Theorem 2.3.
@@ 3. Rees algebra of Almost Complete Intersections
This section discusses the resolutions of the Rees algebra of some classes of ideals $I$ of linear type satisfying a relation between their minimal number of generators and their respective heights. The resolution is obtained as the mapping cone between the resolution of $I$ and the Buchsbaum-Rim complex, corresponding to the generators of $I$. The idea has been discussed in [5] for constructing resolutions of equigenerated ideals. We try to extend the result to include the non-equigenerated case as well. As a consequence of this, minimal bigarded free resolutions of almost complete intersections are obtained.
The following are the building blocks for constructing the minimal bigraded free resolutions given in Theorem 3.1.
@@
    $\epsilon(e_{i})=\sum_{m\subset k,\,|m|=2}\text{sgn}(m\subset k)(\text{det}(\psi_{m}))e_{k\backslash m}$ (8)
@@
    $\sigma_{i}(\beta\otimes\gamma)=\sum_{s,t}[\psi^{*}(\beta_{s}^{\prime})](\gamma_{t}^{\prime})\cdot\gamma_{t}^{\prime\prime}\otimes\beta_{s}^{\prime\prime}.$ (9)
@@
For $i=1,\ldots,l$, we have $\psi^{*}(g_{1}^{*})(e_{i})=f_{i}$ and $\psi^{*}(g_{2}^{*})(e_{i})=y_{i}$. This clearly indicates that the way the differentials are defined means that the degrees of the generators $f_{i}$ are crucial in determining the bishifts in the Buchsbaum-Rim complex.
Let,
.: $\mathbb{C^{\prime}.:}\hskip 85.35826pt\;0\longrightarrow\wedge^{l}F\otimes(\text{Sym}_{l-3}G)^{*}\stackrel{{\scriptstyle\sigma_{l-2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\sigma_{2}}}{{\longrightarrow}}\wedge^{3}{F}\otimes(\text{Sym}_{0}G)^{*}$ (10)
denote the truncated Buchsbaum-Rim complex where $C_{i}^{\prime}=\wedge^{i+2}{F}\otimes(\text{Sym}_{i-1}G)^{*}$, and
$\mathbb{F^{\prime}.:}\hskip 85.35826pt\ldots\stackrel{{\scriptstyle\phi_{4}^{\prime}}}{{\longrightarrow}}\bigoplus_{j\geq 3}S(-j,-1)^{\beta_{3j}}\stackrel{{\scriptstyle\phi_{3}^{\prime}}}{{\longrightarrow}}\bigoplus_{j\geq 2}S(-j,-1)^{\beta_{2j}}$ (11)
denote the truncated complex $\mathbb{F.}\otimes S$ where the differentials $\phi_{i}^{\prime}$ denotes the base change of $\phi_{i}$ to $S$.
Next we mention a map of complexes from $\mathbb{C^{\prime}.}$ to $\mathbb{F^{\prime}.}$ mentioned in [5].
Let $\nu$ be a map of complexes from the Koszul complex corresponding to $\phi_{1}$, denoted by (.$(\wedge^{.}A^{l},c.)$, to $(\mathbb{F.},\phi.)$, the minimal bigraded free resolution of $I$, obtained by extending the identity map at degree 00 and degree $1$ such that the diagram commutes. Further, let $\nu^{\prime}$ denote the base change of $\nu$ to $S$. This map $\nu^{\prime}$ helps in defining the required map of complexes $\alpha:\mathbb{C^{\prime}.}\rightarrow\mathbb{F^{\prime}.}$ which has the following form.
For $21\leq i\leq l-$ 2, $\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\rightarrow F_{i+1}^{\prime}$ is given as the composition of the following two maps.
$\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\stackrel{{\scriptstyle p\otimes\tau}}{{\longrightarrow}}\wedge^{1}F\otimes\wedge^{i+1}F\stackrel{{\scriptstyle\tiny{\begin{bmatrix}y_{1}&\cdots&y_{l}\end{bmatrix}}\otimes\nu_{i+1}^{\prime}}}{{\longrightarrow}}S\otimes F_{i+1}^{\prime}$ (12)
@@
$\cdots\stackrel{{\scriptstyle\delta_{4}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 4}S(-j,-1)^{\beta_{4j}}\\ \oplus\\ C_{2}^{\prime}\end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 3}S(-j,-1)^{\beta_{3j}}\\ \oplus\\ C_{1}^{\prime}\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 2}S(-j,-1)^{\beta_{2j}}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0.$ (13)
@@
$C_{i}^{\prime}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$
are the modules which occur in the Buchsbaum-Rim complex $\mathbb{C^{\prime}}.$ corresponding to the matrix $\begin{bmatrix}f_{1}&\cdots&f_{n+1}\\ y_{1}&\cdots&y_{n+1}\\ \end{bmatrix}$ with differentials denoted by $\sigma_{i}$.
Let $\alpha:\mathbb{C^{\prime}}.\rightarrow\mathbb{F^{\prime}}.$ be the map of complexes given in Equation (12). Then for all $i=1,\ldots,\operatorname{proj\,dim}_{S}\mathcal{R}(I)$, the differentials $\delta_{i}$ are of the following form.
$\delta_{1}=\begin{bmatrix}y_{1}&\ldots&y_{n+1}\end{bmatrix}\circ\phi_{2}^{\prime},\;\;\delta_{2}=\begin{pmatrix}\alpha_{1}&\phi_{3}^{\prime}\end{pmatrix},\;\;\delta_{i}=\begin{pmatrix}\phi_{i+1}^{\prime}&\alpha_{i-1}\\ O&-\sigma_{i-1}\end{pmatrix}.$
@@
Here for $i=1,\ldots,n+1$, $\psi^{*}(g_{1}^{*})(e_{i})=f_{i}$ and $\psi^{*}(g_{2}^{*})(e_{i})=y_{i}$ which determines the entries in the matrices corresponding to the differentials $\sigma_{i}^{\prime}s$. Then, the way the differentials are defined in Equation (9), one obtains
$C_{i}^{\prime}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$ (16)
for $i=2,\ldots,n-1$.
Consider (.$(\wedge^{.}B^{n+1},k.)$ to be the Koszul complex corresponding to $\phi_{1}$. Now since for $\{j_{1},j_{2}\}\subseteq[n+1]$ such that $j_{1}\neq j_{2}$, $k_{2}(e_{j_{1}}\wedge e_{j_{2}})=f_{j_{1}}e_{j_{2}}-f_{j_{2}}e_{j_{1}}\in$ Ker ϕ1=subscriptitalic-$\phi_{1}=$ Im $\phi_{2}$, there exists $\beta_{j_{1}j_{2}}\in F_{2}$ such that $\phi_{2}(\beta_{j_{1}j_{2}})=i(k_{2}(e_{j_{1}}\wedge e_{j_{2}}))$ where $i$ is the natural map from $\wedge^{0}B^{n+1}$ to $F_{1}$. Let $\alpha:\mathbb{C^{\prime}.}\rightarrow\mathbb{F^{\prime}.}$ be the map of complexes given in Equation (12) where $\mathbb{C^{\prime}.}$ and $\mathbb{F^{\prime}.}$ are the complexes given in Equations (10) and (11) respectively. Then $\alpha_{1}:\wedge^{3}{S^{n+1}}\otimes(\operatorname{Sym}_{0}S^{2})^{*}\rightarrow F_{2}^{\prime}$ will be defined as, for $\{j_{1},j_{2},j_{3}\}\subseteq[n+1]$,
$\alpha_{1}((e_{j_{1}}\wedge e_{j_{2}}\wedge e_{j_{3}})\otimes 1)=y_{j_{1}}\beta_{j_{2}j_{3}}-y_{j_{2}}\beta_{j_{1}j_{3}}+y_{j_{3}}\beta_{j_{1}j_{2}}.$
Thus, from the way the map $\alpha_{1}$ is defined, one obtains,
$C_{1}^{\prime}=\bigoplus_{1\leq j_{1}<j_{2}<j_{3}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+d_{j_{3}}),-2).$ (17)
With the details above, one obtains the required resolution (13) by considering the mapping cone of $\alpha$ and following arguments similar to those in [5]. From the construction of the Complex (13), it is evident that the bishifts in the resolution of the Rees algebra of $I$ is derived from the information in the Complexes (14) and (15) and Equations (16) and (17). Further, the resolution is minimal since the entries in the matrices given by the differentials $\delta_{i}$ are from $\mathfrak{m}S$, where $\mathfrak{m}$ is the bigraded maximal ideal of $S$. Note that the linear shifts in the second variable are contributed by the differentials in the minimal free resolution of $B/I$, while the higher degree shifts are contributed by the $\alpha_{i}^{\prime}s$ and $\sigma_{i}^{\prime}s$. ∎
@@
Note that in Theorem 3.1 if $I$ is considered to be a graded ideal of $B$ with $\mu(I)=\text{ht}(I)+1$ and not necessarily of linear type, then the Complex (13) gives a minimal bigraded free resolution of the $\operatorname{Sym}(I)$.
@@
The following observations can be made on the nature of the generators of $I$ as a consequence of the explicit minimal bigraded free resolutions of the symmetric algebra of $I$ obtained from Theorem 3.1.
Note that for an ideal $I$ in $B$, its generating set $\{a_{1},\ldots,a_{n}\}$ forms a strong $s$-sequence (w.r.t an admissible term order for the monomials in $y_{i}$ with $y_{1}<y_{2}<\ldots<y_{n}$) if for $\operatorname{Sym}(I)\cong S/J$ where $S$ is a bigraded polynomial ring $S=B[y_{1},\ldots,y_{n}]$ and $J$ the defining ideal of the symmetric algebra, $in(J)=\langle L_{1}y_{1},\ldots,L_{n}y_{n}\rangle$ where $L_{i}=(a_{1},\ldots,a_{i-1}:a_{i})$ and $L_{1}\subseteq L_{2}\subseteq\ldots\subseteq L_{n}$ (cf. [17]).
@@
From the minimal bigraded free resolution of the $\operatorname{Sym}(I)$ in Complex (13), it is clear that the $y$-regularity of the $\operatorname{Sym}(I)$ is 00. The result then follows from [30, Corollary 3.2(i)].
Further when the ideal is of linear type, from Complex (13), $y$-regularity of the $\mathcal{R}(I)$ vanishes. Then $I$ is generated by a $d$-sequence follows from [30, Corollary 3.2(ii)]. ∎
@@
    $d$-sequence ideals satisfying the condition that the grade of the ideal is one more than the minimal number of generators of the ideal. In particular, this class includes almost complete intersections.
In the remaining part of this section, we give the explicit resolutions of the Rees algebra of some classes of ideals satisfying the hypothesis of Theorem 3.1, in particular almost complete intersections of grade $2$, grade $3$ etc.